Calibrations Status

Speaker : G. Van Buren ( BNL )


Talk time : 12:15, Duration : 00:15
Run 12
  • dE/dx nSigma issue seen in both pp200 and pp500 (RT ticket 2465)
    • Most prominent in V0 analysis identification of pions and protons
  • CuAu remains uncalibrated


Run 13
  • BeamLine
    • Preliminary calibration in mid-March
    • Mid-fill beam shifts introduced to lower luminosities may be an issue (study ongoing)
  • TPC
    • SpaceCharge & GridLeak:
      • preliminary from Run 12 pp510 in use for Run 13
      • initial calibration attempts using Run 13 data hit some hurdles, then stalled
      • additional R&D in progress:
        • charge in the TPC as a measure of the relevant luminosity
        • usage of GMT (software development is late, but showing progress)
    • Alignment:
      • Cosmic data taken before the Run in FF, and during the Run in RFF
      • Studies ongoing using survey + lasers
      • Hopes to use GMT; GMT will need aligned too
    • Drift Velocity:
      • Automatic calibration method in use for the past few years has struggled due to weak laser signals (seems related to gas chemistry)
      • Modified to use reconstructed central membrane lasers positions as a fall-back method
      • Database will need back-filled
    • T0 & Twist:
      • Attempting to use recent low luminosity data (new workforce in training)
  • EMC2
    • Trigger pedestals: ongoing discussion of what gets stored in the database (may introduce new tables)
  • VPD & TOF
  • PXL
    • Engineering Run - won't be followed for physics production readiness